OpenStudy (anonymous):

how do you find the u*v and the angle between u and v to the nearest degree? u=2i+j, v=3i-2j Please show me how!

OpenStudy (zarkon):

\[(a,b)\cdot(c,d)=ac+bd\]

OpenStudy (zarkon):

\[u\cdot v=(2i+j)\cdot(3i-2j)=2\times3+1\times(-2)\]

OpenStudy (zarkon):

\[\theta=\cos^{-1}\left(\frac{u\cdot v}{\|u\|\|v\|}\right)\]